OpenStudy (anonymous):

A sample of gas has a volume of 8.00 L at 20.0 ° C and 700. torr. What will be its volume at STP?

OpenStudy (chmvijay):

use the equation PV=n RT

OpenStudy (chmvijay):

P1V1 /T1 =P2V2/T2 can you solve now :) for STP you take For STP, use 273K(T1), 760 torr(P1), V1=? P2 =700 torr ,T2=20+273 =293 K ,V2=8L

OpenStudy (anonymous):

i have tried that equation a few times and still get teh wrong answer

OpenStudy (chmvijay):

what is the answer you are getting

OpenStudy (chmvijay):

LOL tell me quickly :)

OpenStudy (anonymous):

ok ..sorry, trying to finish a timed quiz..i will tell you

OpenStudy (anonymous):

so i do (700x8.00)/293.15=(760)v2/273.15

OpenStudy (anonymous):

I get 2.78 on right, multiply to cancel on right then multiply on left..i get...

OpenStudy (anonymous):

53.15 as v2

OpenStudy (anonymous):

that's not right

OpenStudy (chmvijay):

LOL u get 6.82 Litres

OpenStudy (anonymous):

how?

OpenStudy (chmvijay):

760 * V1 /273 =700 *8 /293 V1 =700*8*273 / 760*293 =6.86 L

OpenStudy (anonymous):

i'm looking for v1 then not v2?

OpenStudy (chmvijay):

yaa LOL P1 and V1 are at STP values then V1 value u are looking for :)

OpenStudy (anonymous):

how would i have known that the stp is your number 1 values instead of what was given in the question?

OpenStudy (chmvijay):

sorry that i didn't mentioned it 1 values are for STP i thought by putting those standard values u will understand but u didnt :)

OpenStudy (anonymous):

I worked it out...i got the answer but still a little confused... do you have time for one more?

OpenStudy (chmvijay):

what confusion ???

OpenStudy (anonymous):

this whole chapter is confusing!!

OpenStudy (chmvijay):

LOL :)

OpenStudy (anonymous):

can i ask you one more...how many moles are present in 10.0L of nitrogen gas at STP?

OpenStudy (anonymous):

I know it's easy but i'm definitely not remembering soemthing

OpenStudy (chmvijay):

22.4 litres of any gas = 1 mole 10 litres = 1*10 /22.4 =0.44 mole :)

OpenStudy (anonymous):

how am i canceling the 10.0L of NO?

OpenStudy (anonymous):

so i got as far as 10.0LNO(

OpenStudy (chmvijay):

iam not getting you ??

OpenStudy (anonymous):

ok..so i know this is dimensional analysis..we need to cancel liters

OpenStudy (chmvijay):

yaaaa :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Questions marks are where i'm confused. I can't figure how to cancel the liters and get to moles

OpenStudy (chmvijay):



OpenStudy (anonymous):

ooooooohhhhh...ugh!! i'm such an idiot! THANK YOU!!

OpenStudy (chmvijay):

:)

OpenStudy (anonymous):

Final is tomorrow...i'm overwhelmed and barely understand this...

OpenStudy (chmvijay):

ok all the best :)

OpenStudy (anonymous):

yea..thanx